If Z = 2x-y/ x+y

Find dz/dx and dz/dy


1
Expert's answer
2021-02-03T00:04:23-0500

For the function "z = \frac{2x-y}{x+y}" let us find "\frac{\partial z}{\partial x}" and "\frac{\partial z}{\partial y}:"


"\frac{\partial z}{\partial x}=\frac{2(x+y)-(2x-y)}{(x+y)^2}=\frac{3y}{(x+y)^2}"


"\frac{\partial z}{\partial y}=\frac{-(x+y)-(2x-y)}{(x+y)^2}=\frac{-3x}{(x+y)^2}"